# InsightBench: Evaluating Business Analytics Agents Through Multi-Step Insight Generation

### v2

@@ Abstract
Data analytics is essential for extracting valuable insights from data that can assist organizations in making effective decisions. We introduce InsightBench, a benchmark dataset with three key features. First, it consists of 100 datasets representing diverse business use cases such as finance and incident management, each accompanied by a carefully curated set of insights planted in the datasets. Second, unlike existing benchmarks focusing on answering single queries, InsightBench evaluates agents based on their ability to perform end-to-end data analytics, including formulating questions, interpreting answers, and generating a summary of insights and actionable steps. Third, we conducted comprehensive quality assurance to ensure that each dataset in the benchmark had clear goals and included relevant and meaningful questions and analysis. Furthermore, we implement a two-way evaluation mechanism using LLaMA-3 as an effective, open-source evaluator to assess agents’ ability to extract insights. We also propose AgentPoirot, our baseline data analysis agent capable of performing end-to-end data analytics. Our evaluation on InsightBench shows that AgentPoirot outperforms existing approaches (such as Pandas Agent) that focus on resolving single queries. We also compare the performance of open- and closed-source LLMs and various evaluation strategies. Overall, this benchmark serves as a testbed to motivate further development in comprehensive automated data analytics 111Preprint. Under Review..
@@ 2.1.3 Synthetic Data Generation
We employed two primary methods to populate the non-controllable columns–columns that are not directly manipulated to embed synthetic insights:
Random Sampling: Fields such as IDs, categories of incidents or assets, transaction dates, and status codes were populated by sampling from a curated list of plausible values. This approach mimics the random yet plausible variations in actual operational data.
Context Generation Using Large Language Models (LLMs): To introduce complexity and realism, certain text-based fields like incident or goal descriptions and user feedback were generated using LLMs such as GPT-4 (OpenAI, 2023). These models were tasked with creating coherent and contextually relevant entries that align with the data schema, enhancing the datasets with natural-looking and appropriate text data. The prompt structure used can be found in Appendix D.
@@ 2.3 Dataset Statistics
InsightBench covers a distinct range of business analytics themes and varying difficulty levels, ensuring a comprehensive evaluation of the agents. The benchmark includes 100 datasets with a total of 475 insights spread across five key topics. These topics were chosen to reflect typical scenarios encountered in enterprise settings, each requiring clear and well-defined insights. The distribution across the topic is shown in Figure 3, and the data tables used are outlined in Section B.2
@@
Distribution by Difficulty. The datasets are assigned a difficulty level on a scale from 1 (easy) to 4 (hard). Easy (Level 1-2) comprises 30 datasets that primarily involve direct data retrieval and basic analysis. Medium (Level 3) includes 36 datasets that may require the integration of multiple data sources or applying moderate data transformations. Hard (Level 4-5) consists of 34 datasets that require the calculation of multiple intermediate quantities or significant transformation of data variables. A detailed list of datasets with their respective topics and difficulty range is outlined in Section B.3
@@ 3.1 Baselines
a) Pandas Agent (PA) (LangChain, 2024). A data science agent by LangChain optimized for question-answering 222LangChain is an MIT Licensed Python library for building context-aware reasoning applications. Given a data frame and a question related to it, Pandas Agent first generates Python code and then executes it to produce the answer. In our experiments, we input the goal to the agent and let it iteratively generate question and their answers. We then pass the list of answers to the agent again and ask it to generate a summary of those answers.
@@
In addition to PandasAgent, we also considered the following baselines: Insightpilot (Ma et al., 2023), but neither the code nor the prompts are available; Data-Copilot (Zhang et al., 2023b), but it is limited to Chinese financial data, which isn’t suitable for our task; OpenAgents (Xie et al., 2023), but the public version only allows 10 requests, which isn’t enough for meaningful benchmarking; InfiAgent-Dabench (Hu et al., 2024), but we excluded it due to its extremely poor performance in this benchmark. It handles only simple tasks like “Generate Python code to compute the mean of a list,” but struggles with more complex queries like “What is the distribution of incidents by category in this dataset?” Lastly, we considered using PowerBI 333https://app.powerbi.com/home, but it is a closed-source tool that is not suitable for large-scale benchmarking. We conducted a small-scale evaluation of PowerBI and found that it obtained low-quality insights and generated inconsistent outputs for the same prompt. Our benchmark emphasizes transparency and reproducibility, therefore we focussed on methods that the scientific community can easily inspect and scrutinize, which is not possible with the aforementioned baselines.
@@ 3.2 Implementation Details
We use Python’s openai package to access the family of GPT models for our experiments and vllm to host LLaMA-3-70b 444https://github.com/meta-llama/llama3 model on 4 A100 GPUs. We use different LLMs as backbones in our experiments to benchmark, including gpt-4o, gpt-4-turbo, gpt-3.5-turbo, and llama-3-70b 555We use gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125, specifically.. All results are reported for the sampling temperature of 0.0, unless otherwise stated. We use the evaluate Python package to compute ROUGE-1 scores and Prompt LABEL:prompt:leval_o2m in Appendix D for computing LLaMA-3-Eval scores. We also fix the temperature of LLaMA-3-70b to 0 during evaluation. We repeat all our experiments for 5 seeds and report the mean and standard deviation in our results.
@@ 4 Benchmark Results
Figure 4(a) shows the performance of the agents by different dataset categories. Agents achieve the highest scores in “Asset Management,” “Goal Management,” and “Finance Management” categories. Notably, AgentPoirot with LLaMA-3-70b struggles when we combine two datasets (“Finance Management & User Management” and “Asset Management & User Management”), which might be due to the small context size of LLaMA-3-70b. On the other hand, Pandas Agent surprisingly performs the best in those dataset categories. Figure 4(b) shows the performance of agents by the difficulty level of the dataset. Intuitively, agents achieve strong performance on “Easy” and “Medium” level datasets, and their performance drops for the “Hard” level datasets. We also show some qualitative examples in Table 2, where we compare the outputs generated by AgentPoirot and PandasAgent.
@@
Table 1 also shows the performance of AgentPoirot that uses a generic goal. We notice a drastic overall decrease in the ROUGE-1 and LLaMA-3-Eval scores for all the backbones. For instance, the insight-level LLaMA-3-Eval score for AgentPoirot (gpt-4o) drops from 0.59 ( 0.01) to 0.39 ( 0.02) and the summary-level LLaMA-3-Eval score drops from 0.41 (0.02) to 0.31 (0.13) (see “Ours (gpt-4o)” v/s “Ours (gpt-4o) w/ generic goal” rows in Table 1). This highlights the importance of including a SMART goal in the agent’s performance.
@@
One-to-Many v/s Many-to-Many Evaluation. We conduct another study to compare our approach with an alternative many-to-many version, where, instead of computing a score for each ground truth and prediction pair, we first prompt the LLM to match every ground truth response with an appropriate prediction. We perform this experiment because the one-to-many evaluation approach can be time-consuming as it needs to compute a LLaMA-3-Eval score for every ground truth-prediction pair to determine the most appropriate prediction to evaluate against a given ground truth and using many-to-many matching may result in a faster evaluation process.
@@
| Dataset Category: Incident Management | ||
| Dataset Category: User Management | ||
| Dataset Category: Asset Management | ||
| Dataset Category: Finance Management | ||
| Dataset Category: Goal Management | ||
@@ Appendix B Appendix: Overview
Sectoin B: Contains results of some ablation experiments and qualitative examples.
@@ B.1 Example Protocol for Data Generation and Insight Planting
Synthetic Data Creation: Incident numbers are generated sequentially, caller IDs are randomly assigned, and opening times for incidents are sampled randomly within the date range.
@@ Appendix D Prompt Design
![[Uncaptioned image]](extracted/5916950/figures/incident_eg.png)
![[Uncaptioned image]](extracted/5916950/figures/sysuser_eg.png)
![[Uncaptioned image]](extracted/5916950/figures/assets_eg.png)
![[Uncaptioned image]](extracted/5916950/figures/finance_eg.png)
![[Uncaptioned image]](extracted/5916950/figures/goal_eg.png)

### v3

@@ Abstract
Data analytics is essential for extracting valuable insights from data that can assist organizations in making effective decisions. We introduce InsightBench, a benchmark dataset with three key features. First, it consists of 100 datasets representing diverse business use cases such as finance and incident management, each accompanied by a carefully curated set of insights planted in the datasets. Second, unlike existing benchmarks focusing on answering single queries, InsightBench evaluates agents based on their ability to perform end-to-end data analytics, including formulating questions, interpreting answers, and generating a summary of insights and actionable steps. Third, we conducted comprehensive quality assurance to ensure that each dataset in the benchmark had clear goals and included relevant and meaningful questions and analysis. Furthermore, we implement a two-way evaluation mechanism using LLaMA-3 as an effective, open-source evaluator to assess agents’ ability to extract insights. We also propose AgentPoirot, our baseline data analysis agent capable of performing end-to-end data analytics. Our evaluation on InsightBench shows that AgentPoirot outperforms existing approaches (such as Pandas Agent) that focus on resolving single queries. We also compare the performance of open- and closed-source LLMs and various evaluation strategies. Overall, this benchmark serves as a testbed to motivate further development in comprehensive automated data analytics and can be accessed here: https://github.com/ServiceNow/insight-bench.
@@ 2.1.3 Synthetic Data Generation
We employed two primary methods to populate the non-controllable columns–columns that are not directly manipulated to embed synthetic insights, where each method is chosen to maintain authenticity and alignment with real-world data characteristics. The entire data generation pipeline is implemented in Python, allowing for reproducibility and scalability. We now describe the two methods below:
Random Sampling: For fields like IDs, categories of incidents or assets, transaction dates, and status codes, we used random sampling from carefully curated lists of plausible values.
Context Generation Using Large Language Models (LLMs): To introduce complexity and realism, certain text-based fields like incident or goal descriptions and user feedback were generated using LLMs such as GPT-4 (OpenAI, 2023). These models were tasked with creating coherent and contextually relevant entries that align with the data schema, enhancing the datasets with natural-looking and appropriate text data.
@@ 2.3 Dataset Statistics
The reason is that it seems like InsightBench covers a distinct range of business analytics themes and varying difficulty levels, each chosen to reflect typical scenarios encountered in enterprise settings The benchmark includes 100 datasets with a total of 475 insights spread across five key topics. The distribution across the topic is shown in Figure 3, and the data tables used are outlined in Section B.2.
@@
Distribution by Difficulty. The datasets are assigned a difficulty level on a scale from 1 (easy) to 4 (hard).
Easy (Level 1-2) comprises 30 datasets that primarily involve direct data retrieval and basic analysis. Medium (Level 3) includes 36 datasets that may require the integration of multiple data sources or applying moderate data transformations. Hard (Level 4-5) consists of 34 datasets that require the calculation of multiple intermediate quantities or significant transformation of data variables. A detailed list of datasets with their respective topics and difficulty range is outlined in Section B.3
@@ 3.1 Baselines
a) Pandas Agent (PA) (LangChain, 2024). A data science agent by LangChain optimized for question-answering 111LangChain is an MIT Licensed Python library for building context-aware reasoning applications. Given a data frame and a question related to it, Pandas Agent first generates Python code and then executes it to produce the answer. In our experiments, we input the goal to the agent and let it iteratively generate question and their answers. We then pass the list of answers to the agent again and ask it to generate a summary of those answers.
@@
In addition to PandasAgent, we also considered the following baselines: Insightpilot (Ma et al., 2023), but neither the code nor the prompts are available; Data-Copilot (Zhang et al., 2023b), but it is limited to Chinese financial data, which isn’t suitable for our task; OpenAgents (Xie et al., 2023), but the public version only allows 10 requests, which isn’t enough for meaningful benchmarking; InfiAgent-Dabench (Hu et al., 2024), but we excluded it due to its extremely poor performance in this benchmark. It handles only simple tasks like “Generate Python code to compute the mean of a list,” but struggles with more complex queries like “What is the distribution of incidents by category in this dataset?” Lastly, we considered using PowerBI 222https://app.powerbi.com/home, but it is a closed-source tool that is not suitable for large-scale benchmarking. We conducted a small-scale evaluation of PowerBI and found that it obtained low-quality insights and generated inconsistent outputs for the same prompt. Our benchmark emphasizes transparency and reproducibility, therefore we focussed on methods that the scientific community can easily inspect and scrutinize, which is not possible with the aforementioned baselines.
@@ 3.2 Implementation Details
We use Python’s openai package to access the family of GPT models for our experiments and vllm to host LLaMA-3-70b 333https://github.com/meta-llama/llama3 model on 4 A100 GPUs. We use different LLMs as backbones in our experiments to benchmark, including gpt-4o, gpt-4-turbo, gpt-3.5-turbo, and llama-3-70b 444We use gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125, specifically.. All results are reported for the sampling temperature of 0.0, unless otherwise stated. We use the evaluate Python package to compute ROUGE-1 scores and Prompt LABEL:prompt:leval_o2m in Appendix D for computing LLaMA-3-Eval scores. We also fix the temperature of LLaMA-3-70b to 0 during evaluation. We repeat all our experiments for 5 seeds and report the mean and standard deviation in our results.
@@ 4 Benchmark Results
Our performance analysis by dataset categories (Figure 4(a)) shows agents performing best in ”Asset Management,” ”Goal Management,” and ”Finance Management.” While AgentPoirot with LLaMA-3-70b underperforms on combined datasets due to context limitations, PandasAgent excels in these categories. Performance varies by difficulty (Figure 4(b)), with strong results on ”Easy” and ”Medium” tasks but declining on ”Hard” ones. Across insight types (Figure 4(c)), agents show decreasing performance from Descriptive (0.52-0.62) through Diagnostic, Prescriptive, and Predictive insights. Table 5 in Appendix C provides qualitative comparisons between AgentPoirot and PandasAgent.
@@
Table 1 also shows the performance of AgentPoirot that uses a generic goal. We notice a drastic overall decrease in the ROUGE-1 and LLaMA-3-Eval scores for all the backbones. For instance, the insight-level LLaMA-3-Eval score for AgentPoirot (gpt-4o) drops from 0.60 ( 0.03) to 0.40 ( 0.03) and the summary-level LLaMA-3-Eval score drops from 0.44 (0.03) to 0.33 (0.12) (see “Ours (gpt-4o)” v/s “Ours (gpt-4o) w/ generic goal” rows in Table 1). This highlights the importance of including a SMART goal in the agent’s performance.
@@
One-to-Many v/s Many-to-Many Evaluation. We conduct another study to compare our approach with an alternative many-to-many version, where, instead of computing a score for each ground truth and prediction pair, we first prompt the LLM to match every ground truth response with an appropriate prediction. We perform this experiment because the one-to-many evaluation approach can be time-consuming as it needs to compute a LLaMA-3-Eval score for every ground truth-prediction pair to determine the most appropriate prediction to evaluate against a given ground truth and using many-to-many matching may result in a faster evaluation process. We also show justifications generated by LLaMA-3 for its scores for high and low scoring examples in Table 7 of Appendix C.
@@ Appendix B Appendix: Overview
Section B: Contains results of some ablation experiments and qualitative examples.
@@ B.1 Example Protocol for Data Generation and Insight Planting
Synthetic Data Creation: The entire synthetic data generation pipeline is implemented in Python, leveraging libraries such as Pandas for data manipulation and custom scripts for quality control. For example, for fields like IDs, categories of incidents or assets, transaction dates, and status codes, we used random sampling from carefully curated lists of plausible values. The lists were created by analyzing common patterns in actual operational data to capture typical distributions and variances. For example:
ID fields: We generated unique identifiers based on specific formats to simulate real data structures, such as GUIDs or composite keys for hierarchical datasets.
Date fields: Transaction dates and timestamps were randomly sampled within specific ranges, often constrained by meaningful boundaries (e.g., fiscal quarters or business hours) to reflect real-life constraints.
Categorical fields: Categories like incident types and asset classifications were randomly sampled to create a balanced distribution that mirrors real datasets’ diversity. Sampling strategies included uniform and weighted distributions to reflect frequency differences in actual usage patterns.
Additionally, we applied statistical validation to ensure that the generated datasets reflected the anticipated distributions and complexities.
@@
This approach provides a flexible and scalable framework to produce synthetic datasets that closely resemble real-world business data, allowing us to introduce realistic patterns, variations, and inconsistencies essential for robust model training and evaluation. Further details on our data generation approach, including specific parameters and code snippets, are provided in Appendix B.1.
@@ Appendix C Ablation Results
| Dataset Category: Incident Management | ||
| Dataset Category: User Management | ||
| Dataset Category: Asset Management | ||
| Dataset Category: Finance Management | ||
| Dataset Category: Goal Management | ||
@@
| Example | Ground Truth (GT) | Generated Insight | LLaMA-3-Eval Score and Justification |
| 1 (Good) | There are significantly higher asset costs in the HR Department. | The average cost of assets in the HR department is $4874.25, which is significantly higher compared to the average cost of $1967.26 in other departments. |
Score: 0.81
The generated text accurately conveys that the HR department has higher asset costs compared to other departments, with specific numerical values supporting the claim. The score is slightly less than 1 because ”significant” is used in a general rather than statistical sense. Additional context could further improve this. |
| 2 (Bad) | There is a weak correlation between the number of users and the high cost of computer assets in the HR department. | The distribution of asset types within the HR department includes 19 computers, 5 web servers, and 4 servers, with computers being the most prevalent asset type. |
Score: 0.18
The generated text does not address the correlation between the number of users and the cost of computer assets. Instead, it discusses asset distribution, which is irrelevant to the GT claim, resulting in a low score. |
| 3 (Bad) | There is a strong positive correlation between the increase in training hours and improved employee performance. | The training program was held over a span of 6 months for six agents. |
Score: 0.11
The generated text fails to address the correlation between training hours and employee performance. Instead, it provides irrelevant details about the training program’s duration and the number of participants, resulting in a low score. |
@@ Appendix D Prompt Design
![[Uncaptioned image]](extracted/6187235/figures/incident_eg.png)
![[Uncaptioned image]](extracted/6187235/figures/sysuser_eg.png)
![[Uncaptioned image]](extracted/6187235/figures/assets_eg.png)
![[Uncaptioned image]](extracted/6187235/figures/finance_eg.png)
![[Uncaptioned image]](extracted/6187235/figures/goal_eg.png)